# When Knockoffs fail: diagnosing and fixing non-exchangeability of Knockoffs

### v1

@@ Appendix B Computation time of non-parametric Knockoffs
We have performed benchmarks to evaluate the computation time needed for the proposed Non-parametric Knockoffs algorithm compared to the Gaussian algorithm. We use the simulation setup described in Section 4 and vary the number of variables $p$. For all values of $p$, we use $n=p$ samples. We first consider using a single CPU and then consider using 40 CPUs. For the Gaussian approach, parallelization is done in the Graphical Lasso covariance estimator via the $n_{jobs}$ argument in scikit-learn (Pedregosa et al., 2011). Note that profiling the code when using a single CPU shows that most of the computation time is spent in covariance estimation. For the Non-parametric approach, the 40 CPUs are used to train Lasso models that predict $X_{j}$ from $\mathbf{X_{-j}}$ in parallel.
@@
Figure 9: Non-parametric and Gaussian Knockoffs computation time. We use either one CPU or 40 CPUs and report the computation time for various problem dimensions in log-log scale. Note that using non-parametric Knockoffs on a single CPU yields around a 3×3$\times$3 × speedup across all problem dimensions compared to Gaussian Knockoffs. Using 40 CPUs, a 5×5$\times$5 × speedup is achieved for problems under $1000$ variables, and a 10×10$\times$10 × to 15×15$\times$15 × speedup for problems of up to $2000$ variables.
@@
Note that using non-parametric Knockoffs on a single CPU yields around a 3×3$\times$3 × speedup across all problem dimensions compared to Gaussian Knockoffs. Using 40 CPUs, a 5×5$\times$5 × speedup is achieved for problems under $1000$ variables, and a 10×10$\times$10 × to 15×15$\times$15 × speedup for problems of up to $2000$ variables.

### v2

@@ Appendix C Computation time of nonparametric knockoffs
We have performed benchmarks to evaluate the computation time needed for the proposed nonparametric knockoffs algorithm compared to the Gaussian algorithm. We use the simulation setup described in Section 4 and vary the number of variables $p$. For all values of $p$, we use $n=p$ samples. We first consider using a single CPU and then consider using 40 CPUs. For the Gaussian approach, parallelization is done in the Graphical Lasso covariance estimator via the $n_{jobs}$ argument in scikit-learn (Pedregosa et al.,, 2011). Note that profiling the code when using a single CPU shows that most of the computation time is spent in covariance estimation. For the nonparametric approach, the 40 CPUs are used to train Lasso models that predict $X_{j}$ from $\mathbf{X_{-j}}$ in parallel. The result is shown in Fig. 9
@@
Figure 9: Nonparametric and Gaussian knockoffs computation time. We use either one CPU or 40 CPUs and report the computation time for various problem dimensions in log-log scale. Note that using nonparametric (Parallel) Knockoffs on a single CPU yields around a 3×3$\times$3 × speedup across all problem dimensions compared to Gaussian knockoffs. Using 40 CPUs, a 5×5$\times$5 × speedup is achieved for problems under $1000$ variables, and a 10×10$\times$10 × to 15×15$\times$15 × speedup for problems of up to $2000$ variables.
@@
As expected, the auto-regressive procedure is much slower than both other approaches: computing knockoffs is nearly 10×10$\times$10 × slower than using the parallel procedure on 40 CPUs for 1000 variables. Using parallel nonparametric knockoffs on a single CPU yields around a 3×3$\times$3 × speedup across all problem dimensions compared to Gaussian knockoffs. Using 40 CPUs, a 5×5$\times$5 × speedup is achieved for problems under $1000$ variables, and a 10×10$\times$10 × to 15×15$\times$15 × speedup for problems of up to $2000$ variables.
Appendix D Theoretical analysis of parallel knockoff generation
We start of by showing that the parallel algorithm does not produce theoretically valid knockoffs. We then turn to evaluating the gap in covariance estimation. In a Gaussian setting, to ensure exchangeability, it is necessary to have
$\left(X_{1},\ldots,X_{p},\tilde{X}_{1},\ldots,\tilde{X}_{p}\right)\sim\mathcal{N}\left(\begin{pmatrix}\mu\\ \mu\end{pmatrix}\begin{pmatrix}\Sigma&\Sigma-\mathrm{diag}(\mathbf{s})\\ \Sigma-\mathrm{diag}(\mathbf{s})&\Sigma\end{pmatrix}\right),$
for some $\mathbf{s}$ such that the covariance matrix is still positive semidefinite (see Candès et al.,, 2018). Using Gaussianity we can rewrite $X_{j}$ using $\mathbf{x}_{-j}$ and the residual from another i.i.d. sample $\mathbf{x}^{\prime}$:
$\tilde{X}_{j}$ $:=\mathbb{E}\left[X_{j}\middle|X_{-j}\right]+X^{\prime}_{j}-\mathbb{E}\left[X^{\prime}_{j}\middle|X^{\prime}_{-j}\right]$
$=\left(\mu_{j}+\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\left(X_{-j}-\mu_{-j}\right)\right)+X^{\prime}_{j}-\left(\mu_{j}+\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\left(X^{\prime}_{-j}-\mu_{-j}\right)\right)$
$=X^{\prime}_{j}+\Sigma_{j,-j}\Sigma^{-1}_{-j,-j}\left(X_{-j}-X^{\prime}_{-j}\right).$
From this formulation, we observe that the mean, variance, and covariance between knockoff and original covariate is preserved:
  • $\mathbb{E}\left[\tilde{X}_{j}\right]=\mathbb{E}\left[X^{\prime}_{j}+\Sigma_{j,-j}\Sigma^{-1}_{-j,-j}\left(X_{-j}-X^{\prime}_{-j}\right)\right]=\mu_{j}.$
  • For any $j\in[1,p]$:
    $\mathbb{E}\left[\left(\tilde{X}_{j}-\mu_{j}\right)^{2}\right]$ $=\mathbb{E}\left[\left(X^{\prime}_{j}-\mu_{j}+\Sigma_{j,-j}\Sigma^{-1}_{-j,-j}\left(X_{-j}-X^{\prime}_{-j}\right)\right)^{2}\right]$
    $=\mathbb{E}\left[\left(X^{\prime}_{j}-\mu_{j}\right)^{2}\right]+\Sigma_{j,-j}\Sigma^{-1}_{-j,-j}\mathbb{E}\left[\left(X_{-j}-X^{\prime}_{-j}\right)\left(X_{-j}-X^{\prime}_{-j}\right)^{\top}\right]\Sigma^{-1}_{-j,-j}\Sigma_{-j,j}$
    $\quad+2\mathbb{E}\left[\left(X^{\prime}_{j}-\mu_{j}\right)\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\left(X_{-j}-X^{\prime}_{-j}\right)\right]$
    $=\Sigma_{j,j}+2\Sigma_{j,-j}\Sigma^{-1}_{-j,-j}\Sigma_{-j,-j}\Sigma^{-1}_{-j,-j}\Sigma_{-j,j}-2\Sigma_{j,-j}\Sigma^{-1}_{-j,-j}\Sigma_{-j,j}$
    $=\Sigma_{j,j}.$
  • Without loss of generality – the same computation works for any $j\neq l\in\{1,\ldots,p\}$ – we derive $\operatorname{Cov}(\tilde{X}_{1},X_{2})$:
    $\mathbb{E}\left[\left(\tilde{X}_{1}-\mu_{1}\right)(X_{2}-\mu_{2})\right]$ $=\mathbb{E}\left[\left(X^{\prime}_{1}-\mu_{1}+\Sigma_{1,-1}\Sigma^{-1}_{-1,-1}\left(X_{-1}-X^{\prime}_{-1}\right)\right)(X_{2}-\mu_{2})\right]$
    $=\Sigma_{1,-1}\Sigma^{-1}_{-1,-1}\mathbb{E}\left[\left(X_{-1}-X^{\prime}_{-1}\right)(X_{2}-\mu_{2})\right]$
    $=\Sigma_{1,-1}\Sigma^{-1}_{-1,-1}\Sigma_{-1,2}$
    $=\Sigma_{1,-1}\Sigma^{-1}_{-1,-1}\Sigma_{-1,-1}\left(\mathbf{1},\mathbf{0},\ldots,\mathbf{0}\right)$
    $=\Sigma_{1,2},$
    where in the second-to-last line we have used that $\Sigma_{-1,2}$ is the first column of $\Sigma_{-1,-1}$, and therefore we could rewrite it as $\Sigma_{-1,-1}\left(\mathbf{1},\mathbf{0},\ldots,\mathbf{0}\right)$.
Without loss of generality, and for readability, we take the first and the second covariates of the knockoff and compute the covariance between them to show that it is not $\Sigma_{1,2}$ as it should be. Indeed, with this procedure, we could either add the residual from the same individual to each coordinate of the knockoff or add different and independent ones. We are going to show that in both cases, the covariance is not the desired one:
Residuals from independent samples:
Let $X^{\prime}\overset{\mathrm{i.i.d.}}{\sim}X^{\prime\prime}$, then we have that the covariance between the first and second knockoff is given by:
$\mathbb{E}$ $\left[\left(\tilde{X}_{1}-\mu_{1}\right)\left(\tilde{X}_{2}-\mu_{2}\right)\right]$
$=\mathbb{E}\left[\left(X^{\prime}_{1}-\mu_{1}+\Sigma_{1,-1}\Sigma_{-1,-1}^{-1}\left(X_{-1}-X^{\prime}_{-1}\right)\left(X^{\prime\prime}_{2}-\mu_{2}+\Sigma_{2,-2}\Sigma_{-2,-2}^{-1}\left(X_{-2}-X^{\prime\prime}_{-2}\right)\right)\right)\right]$
$=\mathbb{E}\left[\Sigma_{1,-1}\Sigma_{-1,-1}^{-1}\left(X_{-1}-X^{\prime}_{-1}\right)\Sigma_{2,-2}\Sigma_{-2,-2}^{-1}\left(X_{-2}-X^{\prime\prime}_{-2}\right)\right]\qquad\text{using $X^{\prime}\perp\!\!\!\perp X^{\prime\prime}$}$
$=\Sigma_{1,-1}\Sigma_{-1,-1}^{-1}\mathbb{E}\left[\left(X^{-1}-X^{\prime}_{-1}\right)\left(X_{-2}-X^{\prime\prime}_{-2}\right)^{\top}\right]\Sigma_{-2,-2}^{-1}\Sigma_{-2,2}\qquad\text{using $a^{\top}=a$ for $a\in\mathbb{R}$}$
$=\Sigma_{1,-1}\Sigma_{-1,-1}^{-1}\Sigma_{-1,-2}\Sigma_{-2,-2}^{-1}\Sigma_{-2,2}.$
To see that this does not coincide with $\Sigma_{1,2}$, we may observe the simple two-dimensional unit variance random variable with covariance $\rho$. In this case, we have that $\Sigma_{1,-1}=\rho$, $\Sigma_{-1,-1}^{-1}=1$, $\Sigma_{-1,-2}=\rho$, $\Sigma^{-1}_{-2,-2}=1$ and $\Sigma_{-2,2}=\rho$, so
$\mathrm{Cov}\left(\tilde{X}^{1},\tilde{X}^{2}\right)=\rho^{3}.$
Residuals from the same sample:
When the knockoffs are constructed using the residuals from the same sample, the covariance is given by:
$\mathbb{E}$ $\left[\left(\tilde{X}_{1}-\mu_{1}\right)\left(\tilde{X}_{2}-\mu_{2}\right)\right]$
$=\mathbb{E}\left[\left(X^{\prime}_{1}-\mu_{1}+\Sigma_{1,-1}\Sigma_{-1,-1}^{-1}\left(X_{-1}-X^{\prime}_{-1}\right)\left(X^{\prime}_{2}-\mu_{2}+\Sigma_{2,-2}\Sigma_{-2,-2}^{-1}\left(X_{-2}-X^{\prime}_{-2}\right)\right)\right)\right]$
$=\mathbb{E}\left[\left(X^{\prime}_{1}-\mu_{1}\right)\left(X^{\prime}_{2}-\mu_{2}\right)\right]+\mathbb{E}\left[\Sigma_{2,-2}\Sigma_{-2,-2}^{-1}\left(X_{-2}-X^{\prime}_{-2}\right)\left(X^{\prime}_{1}-\mu_{1}\right)\right]$
$\qquad+\Sigma_{1,-1}\Sigma_{-1,-1}^{-1}\mathbb{E}\left[\left(X_{-1}-X^{\prime}_{-1}\right)\left(X^{\prime}_{2}-\mu_{2}\right)\right]$
$\qquad+\mathbb{E}\left[\Sigma_{1,-1}\Sigma_{-1,-1}^{-1}\left(X_{-1}-X^{\prime}_{-1}\right)\Sigma_{2,-2}\Sigma^{-1}_{-2,-2}\left(X_{-2}-X^{\prime}_{-2}\right)\right].$
The first term gives $\Sigma_{1,2}$. The second term can be simplified as
$\mathbb{E}\left[\Sigma_{2,-2}\Sigma_{-2,-2}^{-1}\left(X_{-2}-X^{\prime}_{-2}\right)\left(X^{\prime}_{1}-\mu_{1}\right)\right]$ $=-\Sigma_{2,-2}\Sigma_{-2,-2}^{-1}\Sigma_{-2,1}$
$=-\Sigma_{2,-2}\Sigma_{-2,-2}^{-1}\Sigma_{-2,-2}\left(\mathbf{1},\mathbf{0},\ldots,\mathbf{0}\right)=-\Sigma_{2,1}.$
Similarly, the third term provides
$\Sigma_{1,-1}\Sigma_{-1,-1}^{-1}\mathbb{E}\left[\left(X_{-1}-X^{\prime}_{-1}\right)\left(X^{\prime}_{2}-\mu_{2}\right)\right]$ $=-\Sigma_{1,-1}\Sigma_{-1,-1}^{-1}\Sigma_{-1,2}$
$=-\Sigma_{1,-1}\Sigma_{-1,-1}^{-1}\Sigma_{-1,-1}\left(\mathbf{1},\mathbf{0},\ldots,\mathbf{0}\right)=-\Sigma_{1,2}.$
Finally, the last term can be simplified as
$\mathbb{E}$ $\left[\Sigma_{1,-1}\Sigma_{-1,-1}^{-1}\left(X_{-1}-X^{\prime}_{-1}\right)\Sigma_{2,-2}\Sigma^{-1}_{-2,-2}\left(X_{-2}-X^{\prime}_{-2}\right)\right]$
$=\Sigma_{1,-1}\Sigma_{-1,-1}^{-1}\mathbb{E}\left[\left(\mathbf{x}_{-1}-X^{\prime}_{-1}\right)\left(X_{-2}-X^{\prime}_{-2}\right)^{\top}\right]\Sigma_{-2,-2}^{-1}\Sigma_{-2,2}$
$=2\Sigma_{1,-1}\Sigma_{-1,-1}^{-1}\Sigma_{-1,-2}\Sigma_{-2,-2}^{-1}\Sigma_{-2,2}.$
Therefore, combining the previous, we obtain
$\mathbb{E}$ $\left[\left(\tilde{X}_{1}-\mu_{1}\right)\left(\tilde{X}_{2}-\mu_{2}\right)\right]=-\Sigma_{1,2}+2\Sigma_{1,-1}\Sigma_{-1,-1}^{-1}\Sigma_{-1,-2}\Sigma_{-2,-2}^{-1}\Sigma_{-2,2}.$
Similarly as in the independent residuals setting, we can take the two dimension unit variance random variable to see that this will not coincide in general with the covariance $\rho$. Indeed, in this case,
$\mathrm{Cov}(\tilde{X}_{1},\tilde{X}_{2})=-\rho+2\rho^{3}.$
In this example, for both independent and dependent residuals, although the correlation between the knockoff variables are close to the original one when the correlation is either small or large, they do not coincide; therefore, their parallel knockoffs are not theoretically valid.
In Figure 10, we use the setup of section 5 and explore a range of correlation levels by varying the kernel width parameter. The approximation error is practically zero when the data presents strong correlation or none at all. The error is more substantial in intermediate cases.
Figure 10: Covariance estimation error of Parallel generation. We compute the difference between the correct covariance term $\Sigma_{1,2}$ and the one produced by the parallel method, i.e. $\Sigma_{1,-1}\Sigma_{-1,-1}^{-1}\Sigma_{-1,-2}\Sigma_{-2,-2}^{-1}\Sigma_{-2,2}$. We vary the kernel width parameter to explore different correlation levels in the data. Notice that $\Sigma_{1,2}$ is correctly estimated in regimes with practically no correlation and in regimes with very high correlation. The approximation is the least accurate for intermediate correlation levels.